Find 2 consecutive even integers such that the sum of their squares is 244.
Answer by edjones(7311)

Let the integers be x & x+2.
x^2+(x+2)^2=244
x^2+x^2+4x+4=244
2x^2+4x-240=0
x^2+2x-120=0
(x+12)(x-10)=0
x=10, x=-12
The integers are 10 & 12 or -10 & -12.
.
Ed